Mara's father is 12 years older than twice her age. If last year he was exactly three times her age, how old are each now?
Answer by ankor@dixie-net.com(22740)

Let M = Mara's age now
Let F = Father's age now
:
Write an equation for each statement:
"Mara's father is 12 years older than twice her age."
F = 2M + 12
:
" If last year he was exactly three times her age,"
F - 1 = 3M
F = 3M + 1
:
how old are each now?
:
Substitute (2M+12) for F in the above equation:
2M + 12 = 3M + 1
12 - 1 = 3M - 2M
M = 11 is Mara's present age
then
2(11) + 12 = 34 yrs is Father's present age
:
:
Check solution in the statement:
"last year he was exactly three times her age,"
34 - 1 = 3(11)